# Unmasking Correlations in Nuclear Cross Sections with Graph Neural Networks

### v1

@@ IV.2 Predictions of unknown cross sections
Predicting nuclear cross sections across the entire chart in a consistent manner and with well-defined uncertainties is particularly challenging due to the lack of experimental data on radioactive isotopes and the limited predictive power of existing models. In Section III we have shown that both VAE and INR networks can successfully compress four-channel cross-sections with full energy index into compact, low-dimensional representations. This approach exhibits excellent encoding performance. This non-trivial result suggests that the “effective” space of nuclear cross sections is low-dimensional indeed.
@@
The predictive ability of GNNs is dependent on the structure of the data, as observed in a supplementary experiment where we randomly sampled the TENDL graph to generate the training and test set (instead of the windowing function). In this setting the overall error across all models was higher by $\approx$ 30%. Additionally, as discussed previously, the structure of the latent space plays an important role in the optimization choice for the GNN. From Table 1 we see that in the case of the INR, while the ARMAConv model performs as expected, the GatedConvmodel performs worse compared to the VAE-encoding. It is important to note that the structure of the ARMAConv layer accounts for higher-order neighborhood information in the graph compared to the traditional polynomial layers used in GatedConv and GCNConv. Since ARMA filters can capture the global graph structure with a small number of learnable parameters ($\approx$ 250k), it can better estimate the “islands” in the INR latent space hence providing excellent predictive ability for the test set as seen in both Fig. 4 and Fig. 5. The GatedConv model does not have this advantage and hence we observe a drop in performance for the INR-encoding.
@@
Most importantly, the learned relationships between cross sections of different types within a single nucleus, or between different nuclei can pave the way to much more ambitious studies. In the short term, short-distance extrapolation of cross sections of the type $\sigma_{\alpha}(N,Z)\rightarrow\sigma_{\alpha}(N+i,Z+j)$ for channel $\alpha$ should be relatively reliable, especially if $i$ and $j$ are small, i.e., $i,j\leq 4$. Such extrapolations would be especially useful in the context of nucleosynthesis simulations where precision matters less than accuracy. In the longer term, it should be possible to exploit these relationships to estimate covariances. Current libraries are severely limited in covariance data and in many cases, these data are unreliable. Furthermore, cross-material covariances, that is, how uncertainties in a cross section of one nucleus are related to another type of cross section in another nucleus, have never been estimated. By providing a unified framework to compute cross sections, our methodology offers a viable path forward to estimate all such covariances and properly propagate uncertainties from nuclear data into applications.
V Conclusion
@@
In this work, we used variational autoencoders (VAEs) and implicit neural representations (INRs) to learn the latent space of nuclear cross sections, specifically the elastic, inelastic, $(n,2n)$ and $(n,\gamma)$ cross sections in the fast regime. Both types of architecture were trained on the synthetic data from the TENDL nuclear dataset. We then used graph neural networks (GNNs) on the latent representations to learn the relationships between these four types of cross sections. We showed that we could predict cross sections in a rather large test set with remarkably good accuracy.

### v2

@@ V.2 Predictions of unknown cross sections
Predicting nuclear cross sections across the entire chart in a consistent manner and with well-defined uncertainties is particularly challenging due to the lack of experimental data on radioactive isotopes and the limited predictive power of existing models. In Section IV we have shown that both VAE and INR networks can successfully compress four-channel cross-sections with full energy index into compact, low-dimensional representations. This approach exhibits excellent encoding performance. This non-trivial result suggests that the “effective” space of nuclear cross sections is low-dimensional indeed.
@@
The predictive ability of GNNs is dependent on the structure of the data, as observed in a supplementary experiment where we randomly sampled the TENDL graph to generate the training and test set (instead of the windowing function). In this setting the overall error across all models was higher by $\approx$ 30%. Additionally, as discussed previously, the structure of the latent space plays an important role in the optimization choice for the GNN. From Table 1 we see that in the case of the INR, while the ARMAConv model performs as expected, the GatedConvmodel performs worse compared to the VAE-encoding. It is important to note that the structure of the ARMAConv layer accounts for higher-order neighborhood information in the graph compared to the traditional polynomial layers used in GatedConv and GCNConv. Since ARMA filters can capture the global graph structure with a small number of learnable parameters ($\approx$ 250k), it can better estimate the “islands” in the INR hypernet latent space hence providing excellent predictive ability for the test set as seen in both Fig. 4 and Fig. 5. The GatedConv model does not have this advantage and hence we observe a drop in performance for the INR-encoding.
@@
Most importantly, the learned relationships between cross sections of different types within a single nucleus, or between different nuclei can pave the way to much more ambitious studies. In the short term, short-distance extrapolation of cross sections of the type $\sigma_{\alpha}(N,Z)\rightarrow\sigma_{\alpha}(N+i,Z+j)$ for channel $\alpha$ should be relatively reliable, especially if $i$ and $j$ are small, i.e., $i,j\leq 4$. Such extrapolations would be especially useful in the context of nucleosynthesis simulations where precision matters less than accuracy. In the longer term, it should be possible to exploit these relationships to estimate covariances. Current libraries are severely limited in covariance data and in many cases, these data are unreliable. Furthermore, cross-material covariances, that is, how uncertainties in a cross section of one nucleus are related to another type of cross section in another nucleus, have never been estimated. By providing a unified framework to compute cross sections, our methodology offers a viable path forward to estimate all such covariances and properly propagate uncertainties from nuclear data into applications. In the next section, we show how the current trained networks can already give us valuable information how uncertainties propagate across the nuclear chart.
V.3 Propagation of Uncertainties
To illustrate the power of our deep-learning model, we have performed the following “experiment”. We consider the $(n,2n)$ cross section in the 75As nucleus, which we note $\sigma_{i}(E)$, where $i$ simply indexes the position of 75As in the TENDL dataset. We generate 100 random realizations of the $\sigma_{i}(E)$, which we denote $\sigma^{(\alpha)}_{i}(E)$, by considering the covariance information listed in [29], which treats the entire evaluation as a large multivariate normal distribution. We use the VAE encoder to transform each $\sigma^{(\alpha)}_{i}(E)$ into a vector $\mathbf{x}_{i}^{(\alpha)}$ of the latent space. In the next step, we use these latent vectors $\mathbf{x}_{i}^{(\alpha)}$ to define new feature matrices. Keeping the notations of Section III.3, we write $\mathcal{G}^{(\alpha)}=(\mathcal{V}^{(\alpha)},\mathcal{E}^{(\alpha})$ the graph made of nodes $\mathcal{V}^{(\alpha)}$ and edges $\mathcal{E}^{(\alpha)}$. The set of all nodes $\mathcal{V}^{(\alpha)}$ depend on the feature matrix $\mathbf{X}^{(\alpha)}$ associated with the variation $\alpha$. Note that the only differences between any $\mathbf{X}^{(\alpha)}$ and $\mathbf{X}^{(\beta)}$ is the single latent vector corresponding to 75As: $\mathbf{x}_{i}^{(\alpha)}\neq\mathbf{x}_{i}^{(\beta)}$. However, the trained GNN is the set of parameters that define the function $\chi$ in Eq. (4). This means that the predictions from the trained GNN for each $\mathcal{G}^{(\alpha)}$ are all different from one another: the network is unchanged but its input (=the graph $\mathcal{G}^{(\alpha)}$) is.
Let us note $\Sigma_{i}(E)$ the cross section predicted by the GNN in nucleus $i$ based on the graph $\mathcal{G}$. Technically, $\Sigma_{i}(E)$ is obtained by running the decoder of the VAE or INR on the latent vector predicted by the GNN for that nucleus $i$. Since the GNN is not perfect, $\Sigma_{i}(E)\neq\sigma_{i}(E)$. Similarly, we note $\Sigma_{i}^{(\alpha)}(E)$ the cross section predicted by the GNN in the same nucleus $i$ but with the input graph $\mathcal{G}^{(\alpha)}$. We can now quantify the difference in GNN predictions from any arbitrary realization $\alpha$ by computing the deviation
$\varepsilon_{i}^{(\alpha)}=||\Sigma_{i}-\Sigma_{i}^{(\alpha)}||^{2}.$ (9)
The total mean deviation across the entire TENDL dataset is then just given by $\tfrac{1}{n}\sum_{i}^{n}\varepsilon_{i}^{(\alpha)}$.
Figure 9: Deviation (9) between the GNN prediction based on the initial set of TENDL cross sections and the one based on variation $\alpha=1$, for each nucleus in the TENDL dataset. The plot is normalized between $\epsilon_{\rm min}$ and $0.1\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error. The inset corresponds to the blue square centered at $Z=33$ and $N=42$.
In Figure 9, we plot the deviation (9) of the GNN for each nucleus in the TENDL dataset for the variation $\alpha=1$; similar results would be obtained for all other variations. In this experiment, the deviation should be thought of as a proxy to visualize the propagation of uncertainties. It shows how changes of one cross section in one nucleus propagate to neighboring nuclei. As visualized in the inset of the figure, changes are localized near the central nucleus 75As and decay rapidly as we go away from this nucleus.
Figure 10: Deviation (9) in 75As between prediction from the initial graph $\mathcal{G}$ and three linear combinations of the first random realization $\mathcal{G}^{(1)}$; see text for details. The difference is plotted in real units (barns).
Interestingly, changes do no propagate in a linear manner through the GNN. In Fig. 10, we plot the deviation (9) by generating three different graphs from the $\alpha=1$ variation cross section. Instead of replacing the latent vector $\mathbf{x}_{i}$ by the full variation $\mathbf{x}_{i}^{(\alpha)}$, we first introduce “damped” variations of the cross section characterized by the parameter $\lambda$, $\sigma_{i}\rightarrow\sigma_{i}^{(\alpha,\lambda)}=\sigma_{i}+\lambda(\sigma_{i}^{(\alpha)}-\sigma_{i})$. We then encode these $\sigma_{i}^{(\alpha,\lambda)}$ cross sections to define the corresponding feature matrix $\mathbf{X}^{(\alpha,\lambda)}$ before running the GNN and decoding its prediction. Figure 10 was obtained for three different values $\lambda=0.1,0.5,1.0$ of the damping factor. We see that the difference with the original prediction corresponding to the full variation ($\lambda=1.0$) is in fact smaller than for the other two cases.. The reason for this behavior is that each GNN learns an unique mapping function to predict the feature vector at a specific node based on the information from the neighboring nodes. Hence, the way that information is propagated depends on the nature of the mapping function, which is not always linear. However, our method still provides crucial insight into the behavior of neighboring nuclei surrounding the perturbation site.
Figure 11: Predictions of the $(n,2n)$ cross section in 74Ge ($\delta Z=-1$), 75As ($\delta Z=0$), and 76Se ($\delta Z=+1$) for the GNN with the initial graph $\mathcal{G}$ (plain lines, full symbols) and the full variation $\mathcal{G}^{(1)}$ (dashed lines, open symbols). Here, $\delta Z$ refers to the difference in proton number relative to 75As. Cross sections are plotted in real units (barns).
Finally, to visualize the concrete impact of these variations on actual cross sections, we plot in Fig. 11 the full $(n,2n)$ cross section as predicted by the GNN, either with the initial graph $\mathcal{G}$ – marked as ’(ini.)’ – or with the $\alpha=1$ variation graph $\mathcal{G}^{(1)}$ – marked as ’(var.)’. The plots are made for the central nucleus where the variation is generated, 75As, as well as for its two nearest neighbors along the proton number axis, 74Ge ($Z=32$ and $N=42$) and 76Se ($Z=34$ and $N=42$).
Running the GNN on such an updated set of cross sections – in our case, the $\mathcal{G}^{(\alpha)}$ for the variations $\alpha$ of the $(n,2n)$ cross section in 75As – takes less than a second on a laptop. It provides an estimate of how changes in one cross section impacts all other cross sections in all other nuclei. Crucially, this information is completely agnostic: whether the changes in that cross section come from new measurements, better theories, better evaluations or a combination of the above, does not matter. Our approach is completely data driven. As mentioned in Section II, it is simply impossible with the current state of nuclear theory and nuclear data evaluations to obtain the same information.
VI Conclusion
@@
In this work, we used variational autoencoders (VAEs) and implicit neural representations (INRs) to learn the latent space of nuclear cross sections, specifically the elastic, inelastic, $(n,2n)$ and $(n,\gamma)$ cross sections in the fast regime. Both types of architecture were trained on the synthetic data from the TENDL nuclear dataset. We then used graph neural networks (GNNs) on the latent representations to learn the relationships between these four types of cross sections. We showed that we could predict cross sections in a rather large test set with remarkably good accuracy. We also outlined how to quantify the impact new measurements or theoretical calculations would have on the GNN predictions.
@@ Acknowledgements
VII Appendix
In this section, we provide supplementary analyses and ablation studies to support the findings presented in the main paper. These experiments are designed to validate the robustness and generalization capabilities of the proposed framework. We begin by investigating the representation learning phase with the VAE. Following this, we present the results of the downstream GNN prediction task, focusing on consistent performance across different test regions and varying window lengths.
VII.1 Train-Test Performance Trends for VAE
To ensure that our model is not overfitting during the representation learning phase, we analyze the evolution of the loss and PSNR for both the training and testing set during the VAE training. These metrics are presented as a function of epochs in Figure 12. The training and testing loss curves show a steady decrease, converging to similar values by the end of training. Also, the PSNR values for both training and testing increase consistently, improving reconstruction quality over time. The consistent convergence of both metrics indicates minimal overfitting and robust generalization of the learned latent representations.
Figure 12: Loss (dashed curves) and PSNR (plain curves) for the training set (red) and testing set (blue) as a function of epoch during VAE training. Note that the y-axis for the loss is on a logarithmic scale.
VII.2 Generalization Analysis in GNN
Performance consistency across test regions
In the downstream prediction task with the GNN, we conduct an analysis to verify the model’s ability to generalize across different regions of the test set. To this end, we perform evaluation on multiple test regions centered on distinct nuclear chart coordinates while keeping the window length fixed (e.g., $\Delta Z=\Delta N=8$). Figure 13 indicates that test PSNR scores across all randomly selected regions are consistent and demonstrates that the model is not biased toward specific regions of the training set. The three curves with the lowest PSNR correspond to regions centered on 214Tl ($Z=81,N=133$), 219At ($Z=85,N=134$) and 239Pa ($Z=91,N=148$). In these three cases, the TENDL library contains overall fewer isotopes of each element; see Fig. 9. As a result, the number of actual nuclei in the test set is much smaller than the 289 nuclei of most other regions. This lack of samples in these regions results in sparse relationships between neighboring nuclei compared to regions with denser neighboring points. This sparsity limits the GNN’s ability to learn comprehensive connections, leading to lower PSNR scores compared to denser test regions.
Figure 13: PSNR values in the test set as a function of epochs for test regions centered at distinct nuclear chart coordinates $[N_{0},Z_{0}]$. All test regions have the same size characterized by $\Delta Z=\Delta N=8$; only the center nucleus changes.
Performance consistency across window length variations
Figure 14 shows the test PSNR scores across different window lengths for the Gated and ARMA approaches. This analysis highlights how the model performs when varying the size of the test region, where increasing the window length reduces the number of training samples available. Such a setup is crucial to evaluate the model’s ability to generalize effectively, as it forces the model to learn and infer from sparse training data. As see in this figure, the model demonstrates robust and consistent performance across a wide range of window lengths, verifying its capability to generalize well even under challenging conditions.
Figure 14: Test PSNR values across varying window sizes for Gated (left) and ARMA (right) approaches, with the center of the test window fixed at $(Z_{0},N_{0})=(42,60)$.
VII.3 Location of latent clusters
In this section, we show a few more examples of the location of clusters of nuclei identified directly from the VAE or INR latent vectors. Figures 15 quantifies the impact of changing the numbers of clusters to $n_{\rm clus}=6$. The main change is the splitting of the cluster of proton-rich nuclei into two subsets and the appearance of a small cluster of very heavy nuclei.
Figure 15: Location of nuclei in the training set with color code based on which cluster the nucleus belongs too. Black horizontal and vertical lines represent magic numbers. The figure corresponds to the INR latent vectors and: $n_{\rm clus}=6$.
Figure 16 is the exact same figure as Fig.8 in the main text, only it was constructed from the spectral clustering of the VAE latent space (with $n_{\rm clus}=4$). We notice that the overall “diagonal” grouping of nuclei remains, although clusters are a little less compact. In addition, the main cluster that surrounds the valley of stability is not split in two at $N=82$ like in Fig. 8 but instead at $N=126$. It is still interesting that neutron magic numbers seem to be identified as markers of clustering in both methods.
Figure 16: Same as Fig.15. The figure corresponds to the VAE latent vectors and: $n_{\rm clus}=4$.
Finally, Fig. 17 illustrates another clustering method based on the KMeans algorithm. It is constructed from the clustering of the INR latent vectors and corresponds to $n_{\rm clus}=6$. Compared to Fig. 15, it illustrates the impact of a different clustering method. The KMeans method is a little more naive than spectral clustering and makes additional hypotheses about the clusters, in particular that they should be spherical. The main difference between the two methods is that KMeans produces “thinner” clusters. However, the diagonal nature of the clusters and the role of magic neutron numbers in identifying clusters remains.
Figure 17: Same as Fig.15. The figure corresponds to the INR latent vectors and is obtained by the KMeans algorithm instead of spectral clustering and with: $n_{\rm clus}=6$.